# Cross-Domain Graph Data Scaling: A Showcase with Diffusion Models

### v1

@@ Abstract
Models for natural language and images benefit from data scaling behavior: the more data fed into the model, the better they perform. This ’better with more’ phenomenon enables the effectiveness of large-scale pre-training on vast amounts of data. However, current graph pre-training methods struggle to scale up data due to heterogeneity across graphs. To achieve effective data scaling, we aim to develop a general model that is able to capture diverse data patterns of graphs and can be utilized to adaptively help the downstream tasks. To this end, we propose UniAug, a universal graph structure augmentor built on a diffusion model. We first pre-train a discrete diffusion model on thousands of graphs across domains to learn the graph structural patterns. In the downstream phase, we provide adaptive enhancement by conducting graph structure augmentation with the help of the pre-trained diffusion model via guided generation. By leveraging the pre-trained diffusion model for structure augmentation, we consistently achieve performance improvements across various downstream tasks in a plug-and-play manner. To the best of our knowledge, this study represents the first demonstration of a data-scaling graph structure augmentor on graphs across domains. Our implementations are publicly available on https://github.com/WenzhuoTang/UniAug.
@@ 1 Introduction
The effectiveness of existing foundation models radford2021learning ; touvron2023llama ; kirillov2023segment heavily relies on the availability of substantial amounts of data, where the relationship manifests as a scaling behavior between model performance and data scale kaplan2020scaling . Consistent performance gain has been observed with the increasing scale of pre-training data in both Natural Language Processing kaplan2020scaling ; hoffmann2022training and Computer Vision abnar2022exploring ; zhai2022scaling domains. This data scaling phenomenon facilitates the development of general models endowed with extensive knowledge and effective data pattern recognition capabilities. In downstream applications, these models are capable of adaptively achieving performance gains across different tasks.
@@
In the context of graphs, the availability of large-scale graph databases rossi2015network ; hu2020OGB ; snapnets enables possible data scaling across datasets and domains. Existing works have demonstrated graph data scaling following two limited settings: in-domain pre-training xia2023molebert ; liu2024rethinking and task-specific selection for pre-training data cao2023pre . During the pre-training process, each graph in the pre-training pool must be validated as in-domain or relevant to the downstream dataset. Given a specific domain or task, the crucial discriminative data patterns are likely confined to a fixed set mao2024revisiting , leaving other potential patterns in diverse graph data distribution as noisy input. In terms of structure, graphs from different domains are particularly composed of distinct patterns milo2002network , making it hard to transfer across domains. For example, considering the building blocks of the graphs, the motifs shared by ecological food webs differ significantly from those shared by genetic networks milo2002network . Therefore, closely aligning the characteristics of the pre-training graphs and the downstream data both in feature and structure is essential for facilitating positive transfer cao2023pre . As a consequence, the necessity of such meticulous data filtering restricts these methods to scaling up graphs effectively, as they can only utilize a small part of the available data. Given the limitation of the graph pre-training methods, a pertinent question emerges: How can we effectively leverage the increasing scale of graph data across domains?
@@
Rather than focusing solely on data patterns specific to particular domains, we aim to develop a model that has a comprehensive understanding of data patterns inherent across various types of graphs. In line with the principles of data scaling, we hypothesize that incorporating a broader range of training datasets can help the model build an effective and universal graph pattern library, avoiding an overemphasis on major data patterns specific to any single dataset mao2024demystifying . To construct such a general-purpose model, we propose to utilize a diffusion model operating only on the structure as the backbone, for the following key reasons. (1) Current supervised and self-supervised methods tend to capture only specific patterns of graph data, with models designed for particular inductive biases mao2024demystifying ; mao2024revisiting ; xu2018powerful . For instance, graph convolutional networks (GCNs) excel in node-level representation learning by emphasizing homophily, whereas graph-level representation learning benefits from expressive GNNs capable of distinguishing complex graph structures. (2) We opt for a structure-only model due to the heterogeneous feature spaces across graphs, which often include missing features or mismatched semantics mao2024graph . For instance, node features yield completely different interpretations in citation networks, where they represent keywords of documents, compared to molecular networks, where they denote properties of atoms. To this end, we pre-train a structure-only diffusion model on thousands of graphs, which serves as the upstream component of our framework.
@@
In the downstream stage, we employ the pre-trained diffusion model as a Universal graph structure Augmentor (UniAug) to enhance the dataset, where diffusion guidance ho2022classifier ; dhariwal2021diffusion ; gruver2024protein is employed to align the generated structure with the downstream requirements. Specifically, we generate synthetic structures with various guidance objectives, and the resulting graphs consist of generated structures and original node features. This data augmentation paradigm strategically circumvents feature heterogeneity and fully utilizes downstream inductive biases by applying carefully designed downstream models to the augmented graphs in a plug-and-play manner. Empirically, we apply UniAug to graphs from diverse domains and consistently observe performance improvement in node classification, link prediction, and graph property prediction. To the best of our knowledge, this study represents the first demonstration of a data-scaling graph structure augmentor on graphs across domains.
@@ 2 Preliminary and Related Work
Learning from unlabeled graphs Graph self-supervised learning (SSL) methods provide examples of pre-training and fine-tuning paradigm hu2019strategies ; hou2022graphmae ; kim2022graph ; you2021graph ; xu2021self . However, these methods benefit from limited data scaling due to feature heterogeneity, structural pattern differences across domains, and varying downstream inductive biases. It is worth mentioning that DCT liu2024data presents a pre-training and then data augmentation pipeline on molecules. Despite its impressive performance improvement on graph-level tasks, DCT is bounded with molecules and thus the use cases are limited.
Graph data augmentation There have been many published works exploring graph data augmentation (GDA) since the introduction of graph neural networks (GNNs), with a focus on node-level park2021metropolis ; liu2022local ; azabou2023half , link-level zhao2022learning ; nguyen2024diffusion , and graph-level han2022g ; ling2023graph ; luo2022automated ; liu2022graph ; kong2022robust . These GDA methods have been generally designed for specific tasks or particular aspects of graph data. In addition, they are often tailored for a single dataset and struggle to transfer to unseen patterns, which limits their generalizability to a broader class of applications.
Diffusion models on graphs Diffusion models ho2020denoising ; song2021scorebased ; rombach2022high are latent variable models that learn data distribution by gradually adding noise into the data and then recovering the clean input. Existing diffusion models on graphs can be classified into two main categories depending on the type of noise injected, i.e. Gaussian or discrete. Previous works employed Gaussian diffusion models both on general graphs niu2020permutation ; jo2022score and molecules shi2021learning ; xu2022geodiff . However, adding Gaussian noise into the adjacency matrix will destroy the sparsity of the graph, which hinders the scalability of the diffusion models haefeli2022diffusion . Recent works adapted discrete diffusion models to graphs with categorical transition kernels vignac2023digress ; chen2023efficient ; chen2023dexplainer . We denote the adjacency matrix of a graph as $\mathbf{A}^{0}\in\{0,1\}^{n\times n}$ with $n$ nodes. With details in Appendix A, we write the forward process to corrupt the adjacency matrix into a sequence of latent variables as Bernoulli distribution
[…]
@@
$p_{\theta}\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t}\right)\propto\sum_{\widetilde{\mathbf{A}}_{0}}q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\widetilde{\mathbf{A}}_{0}\right)\tilde{p}_{\theta}\left(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t}\right),$
where $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t})$ represents the denoising network that predicts the original adjacency matrix from the noisy adjacency matrix. The parameters are estimated by optimizing the variational lower bound on the negative log-likelihood austin2021structured
[…]
@@ 3 Method
In this section, our goal is to build UniAug to understand the diverse structure patterns of graphs and perform data augmentation with a range of objectives. As illustrated in Fig.1, UniAug consists of two main components: a pre-trained diffusion model and the downstream adaptation through structure augmentation. We first collect thousands of graphs from varied domains with diverse patterns. To construct a general model free of downstream inductive biases, we train a self-conditioned discrete diffusion model on graph structures. In the downstream stage, we train an MLP guidance head on top of the diffusion model with objectives across different levels of granularity. We then augment the downstream dataset by generating synthetic structures through guided generation, where the augmented graph is composed of generated structures and original node features. Subsequently, we apply the augmented data to train a task-specific model for performing downstream tasks. Below, we elaborate on the data collection process, the architecture of the discrete diffusion model, and the guidance objectives employed.
@@ 3.1 Pre-training data collection
Figure 2: Normalized structural properties.
In light of the data scaling spirit, we expect our pre-training data to contain diverse data patterns with sufficient volume. As graphs from different domains exhibit different patterns milo2002network , we wish to build a collection of graphs from numerous domains to enable a universal graph pattern library with pre-training. Within the publicly available graph databases, Network Repository rossi2015network provides a comprehensive collection of graphs with varied scales from different domains, such as biological networks, chemical networks, social networks, and many more. Among the thousands of graphs in the Network Repository, some of them contain irregular patterns, including multiple levels of edges, extremely high density, et cetera. To ensure the quality of the graphs, we analyze the graph properties following Xu et al. xu2023better and filter out the outliers. In addition, we observe that the coverage of graphs in the Network Repository is incomplete according to the network entropy and scale-free exponent, as we observe a relatively scattered space in the middle of Fig. 2. To fill in the gap, we include a $1000$ graphs subset of the GitHub Star dataset rozemberczki2020karate . The selected graphs are utilized to train a discrete diffusion model.
@@ 3.2 Pre-training through diffusion model
Diffusion models have demonstrated the ability to facilitate transferability from a data augmentation perspective on the images trabucco2024effective ; you2024diffusion ; he2023is . Unlike the traditional hand-crafted data augmentation methods, diffusion models are capable of producing more diverse patterns with high quality trabucco2024effective . With the aid of diffusion guidance ho2022classifier ; dhariwal2021diffusion , these methods can achieve domain customization tailored to specific semantic spaces you2024diffusion ; he2023is . Despite the success of data augmentation through diffusion models on images, the non-Euclidean nature of graph structures poses challenges for data-centric learning on graphs. In addition, the fact that most graphs in the Network Repository are unlabeled exacerbates the challenges, as the absence of labeled data results in substantially lower generation quality for diffusion models dhariwal2021diffusion ; bao2022conditional .
@@
To address the aforementioned challenges, we propose to construct a self-conditioned discrete diffusion model on graph structures. Unlike Gaussian-based diffusion models, discrete diffusion models hoogeboom2021argmax ; austin2021structured ; campbell2022continuous ; vignac2023digress operate with discrete transition kernels between latent variables, as shown in Section 2. The key reason we opt for the discrete diffusion models lies in the sparse nature of graphs, where adding Gaussian noise into the adjacency matrix will result in a dense graph haefeli2022diffusion . On the contrary, discrete diffusion models effectively preserve the sparse structure of graphs during the diffusion process, thus maintaining the efficiency of the models on graphs.
@@
To accommodate for unlabeled graphs, we adopt a self-supervised labeling strategy as an auxiliary conditioning procedure gao2022large ; hu2023self . By leveraging the self-labeling technique, we are able to upscale the diffusion model to data with more diverse patterns gao2022large . The self-labeling technique requires two components, i.e., a feature extractor and a self-supervised annotator.
@@
Feature extractor. We extract graph-level features by calculating graph properties, including the number of nodes, density, network entropy, average degree, degree variance, and scale-free exponent following Xu et al. xu2023better . The first two represent the scale of the graph corresponding to nodes and edges, and the rest indicate the amount of information contained within a graph xu2023better . We compute the properties of one graph and concatenate them to get a graph-level representation.
@@
Self-supervised annotator. To assign labels to graphs in a self-supervised manner, we employ clustering algorithms on the graph-level representations. The number of clusters is determined jointly by the silhouette score rousseeuw1987silhouettes and the separation of the graphs. The candidates of the number of clusters are chosen to ensure different clusters are well separated. Among the candidates, we select the final number of clusters by maximizing the mean Silhouette Coefficient of all samples.
@@
Next we detail the parameterization of the denoising model $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{A}^{t})$ with the self-assigned graph-level labels $\mathbf{k}$. The denoising model recovers the edges of the original adjacency matrix by predicting the connectivity of the upper triangle, which can be formulated as a link prediction problem zhang2018link ; kumar2020link . Following the link prediction setup, the denoising model is composed of a graph transformer (GT) shi2020masked and an MLP link predictor. Denote the hidden dimension as $d$, we treat the node degrees as node features and utilize a linear mapping $f_{d}:\mathbb{R}\mapsto\mathbb{R}^{d}$ to match the dimension. Similarly, we utilize another linear mapping $f_{t}:\mathbb{R}\mapsto\mathbb{R}^{d}$ for timestep $t$ and learnable embeddings $f_{k}:\{0,\dots,K\}\mapsto\mathbb{R}^{d}$ for labels $\mathbf{k}$, where $K$ is the number of clusters. The outputs are summed together and then fed into the GT. Mathematically, we have
@@
3.3 Downstream adaptation through data augmentation
The downstream phase of UniAug is to augment the graph topology through guided generation. This guidance process serves to provide downstream semantics for the diffusion model, thus bridging the gap between the pre-training distribution and the downstream datasets. Among the techniques for diffusion guidance, gradient-based methods dhariwal2021diffusion ; gruver2024protein offer versatile approaches by incorporating external conditions that are not present during training. For the discrete diffusion process, we opt for the gradient-based NOS method gruver2024protein due to its flexibility and efficiency. Specifically, we build an MLP regression head $g_{\theta}:\mathbb{R}^{d}\mapsto\mathbb{R}^{r}$ that takes the hidden representations $\mathbf{h}^{t}$ as the input and outputs the guidance objective of dimension $r$. Denote $\tau$ as the temperature, $\gamma$ as the step-size, $\lambda$ as the regularization strength, and $\varepsilon$ drawn from $\mathcal{N}(0,I)$, we sample from $\tilde{p}^{\prime}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\propto\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\exp\left(g_{\theta}\left(\mathbf{h}^{t}\right)\right)$ via Langevin dynamics
Node level. Node labels provide the supervision signal for node classification tasks. Beyond node labels, node degrees are a fundamental factor in the evolutionary process of a graph liu2011controllability . From the perspective of network analysis, centrality measures indicate the importance of nodes from various viewpoints borgatti2005centrality . Empirically, we observe that utilizing different node-level heuristics as guidance targets tends to yield similar outcomes. Therefore, we focus on node labels and node degrees.
Edge level. Edge-level heuristics can be broadly classified into two categories: local structural heuristics, such as Common Neighbor and Adamic Adar adamic2003friends , and global structural heuristics, such as Katz katz1953new and SimRank jeh2002simrank . Similar to node-level heuristics, empirical observations suggest that different edge-level heuristics tend to yield comparable guidance effects. In this work, we focus on the Common Neighbors (CN) heuristic due to its efficiency. Another edge-level guidance objective is to recover the adjacency matrix from the node representations in a link prediction way, similar to how we parameterize the denoising network. We anticipate that such link prediction objective helps to align the generated graph with the downstream data on the granularity of edges.
Graph level. Graph labels offer the supervision signal for graph classes or regression targets. In addition, we incorporate graph-level properties xu2023better as quantitative measures to bridge the gap between the pre-training distribution and the downstream dataset. We empirically observe that graph label guidance offers significantly higher performance boosts compared to properties on graph-level tasks. Therefore, we focus on graph labels in our experiments.
We provide our choice of objectives for each task in Appendix B. We note that all the above objectives are natural choices inspired by heuristics and downstream tasks. There exist many other self-supervised objectives to be explored, such as community-level spectral change tan2024community and motif occurrence prediction rong2020self . We leave the study of these objectives as one future work. With the diffusion guidance, we assemble the augmented graphs with generated structures and original node features. The augmented graphs are then fed into downstream-specific GNNs.
3.4 Comparison to existing methods

### v3

@@ Abstract
Models for natural language and images benefit from data scaling behavior: the more data fed into the model, the better they perform. This ’better with more’ phenomenon enables the effectiveness of large-scale pre-training on vast amounts of data. However, current graph pre-training methods struggle to scale up data due to heterogeneity across graphs. To achieve effective data scaling, we aim to develop a general model that is able to capture diverse data patterns of graphs and can be utilized to adaptively help the downstream tasks. To this end, we propose UniAug, a universal graph structure augmentor built on a diffusion model. We first pre-train a discrete diffusion model on thousands of graphs across domains to learn the graph structural patterns. In the downstream phase, we provide adaptive enhancement by conducting graph structure augmentation with the help of the pre-trained diffusion model via guided generation. By leveraging the pre-trained diffusion model for structure augmentation, we consistently achieve performance improvements across various downstream tasks in a plug-and-play manner. To the best of our knowledge, this study represents the first demonstration of a data-scaling graph structure augmentor on graphs across domains.
@@ 1 Introduction
The effectiveness of existing foundation models (radford2021learning, ; touvron2023llama, ; kirillov2023segment, ) heavily relies on the availability of substantial amounts of data, where the relationship manifests as a scaling behavior between model performance and data scale (kaplan2020scaling, ). Consistent performance gain has been observed with the increasing scale of pre-training data in both Natural Language Processing (kaplan2020scaling, ; hoffmann2022training, ) and Computer Vision (abnar2022exploring, ; zhai2022scaling, ) domains. This data scaling phenomenon facilitates the development of general models endowed with extensive knowledge and effective data pattern recognition capabilities. In downstream applications, these models are capable of adaptively achieving performance gains across tasks.
@@
In the context of graphs, the availability of large-scale graph databases (rossi2015network, ; hu2020OGB, ; snapnets, ) enables possible data scaling across datasets and domains. Existing works have demonstrated graph data scaling following two limited settings: in-domain pre-training (xia2023molebert, ; liu2024rethinking, ) and task-specific selection for pre-training data (cao2023pre, ). During the pre-training process, each graph in the pre-training pool must be validated as in-domain or relevant to the downstream dataset. Given a specific domain or task, the crucial discriminative data patterns are likely confined to a fixed set (mao2024revisiting, ), leaving other potential patterns in diverse graph data distribution as noisy input. In terms of structure, graphs from different domains are particularly composed of varied patterns (milo2002network, ), making it hard to transfer across domains. For example, considering the building blocks of the graphs, the motifs shared by the World Wide Web hyperlinks only partially align with those shared by genetic networks (milo2002network, ). Therefore, closely aligning the characteristics of the pre-training graphs and the downstream data both in feature and structure is essential for facilitating positive transfer (cao2023pre, ). As a consequence, the necessity of such meticulous data filtering restricts these methods from scaling up graphs effectively, as they can only utilize a small part of the available data. Given the limitation of the graph pre-training methods, a pertinent question emerges: How can we effectively leverage the increasing scale of graph data across domains?
@@
Rather than focusing solely on data patterns specific to particular domains, we aim to develop a model that has a comprehensive understanding of data patterns inherent across various types of graphs. In line with the principles of data scaling, we hypothesize that incorporating a broader range of training datasets can help the model build an effective and universal graph pattern library, avoiding an overemphasis on major data patterns specific to any single dataset (mao2024demystifying, ). To construct such a general-purpose model, we propose to utilize a diffusion model operating only on the structure as the backbone, for the following key reasons. (1) Unlike features, graph structures follow a uniform construction principle, namely, the connections between nodes. This allows for positive transfer across domains when the upstream and downstream data exhibit similar topological patterns (cao2023pre, ). In particular, while the graph representations of neurons and forward electronic circuits are derived from distinct domains, they still share common motifs (milo2002network, ). (2) Current supervised and self-supervised methods tend to capture only specific patterns of graph data, with models designed for particular inductive biases (mao2024demystifying, ; mao2024revisiting, ; xu2018powerful, ). For instance, graph convolutional networks (GCNs) excel in node-level representation learning by emphasizing homophily, whereas graph-level representation learning benefits from expressive GNNs capable of distinguishing complex graph structures. (3) We opt for a structure-only model due to the heterogeneous feature spaces across graphs, which often include missing features or mismatched semantics (mao2024graph, ). For instance, node features yield completely different interpretations in citation networks, where they represent keywords of documents, compared to molecular networks, where they denote properties of atoms. To this end, we pre-train a structure-only diffusion model on thousands of graphs, which serves as the upstream component of our framework.
@@
In the downstream stage, we employ the pre-trained diffusion model as a Universal graph structure Augmentor (UniAug) to enhance the dataset, where diffusion guidance (ho2022classifier, ; dhariwal2021diffusion, ; gruver2024protein, ) is employed to align the generated structure with the downstream requirements. Specifically, we generate synthetic structures with various guidance objectives, and the resulting graphs consist of generated structures and original node features. This data augmentation paradigm strategically circumvents feature heterogeneity and fully utilizes downstream inductive biases by applying carefully designed downstream models to the augmented graphs in a plug-and-play manner. Empirically, we apply UniAug to graphs from diverse domains and consistently observe performance improvement in node classification, link prediction, and graph property prediction. To the best of our knowledge, this study represents the first demonstration of a cross-domain data-scaling graph structure augmentor.
@@ 2 Preliminary and Related Work
Learning from unlabeled graphs.
Graph self-supervised learning (SSL) methods provide examples of pre-training and fine-tuning paradigm (hu2019strategies, ; hou2022graphmae, ; kim2022graph, ; you2021graph, ; xu2021self, ). However, these methods benefit from limited data scaling due to feature heterogeneity, structural pattern differences across domains, and varying downstream inductive biases. It is worth mentioning that DCT (liu2024data, ) presents a pre-training and then data augmentation pipeline on molecules. Despite its impressive performance improvement on graph-level tasks, DCT is bounded with molecules and thus the use cases are limited.
Graph data augmentation.
There have been many published works exploring graph data augmentation (GDA) since the introduction of graph neural networks (GNNs), with a focus on node-level (park2021metropolis, ; liu2022local, ; azabou2023half, ), link-level (zhao2022learning, ; nguyen2024diffusion, ), and graph-level (han2022g, ; ling2023graph, ; luo2022automated, ; liu2022graph, ; kong2022robust, ). These GDA methods have been generally designed for specific tasks or particular aspects of graph data. In addition, they are often tailored for a single dataset and struggle to transfer to unseen patterns, which limits their generalizability to a broader class of applications.
Diffusion models on graphs.
Diffusion models (ho2020denoising, ; song2021scorebased, ; rombach2022high, ) are latent variable models that learn data distribution by gradually adding noise into the data and then recovering the clean input. Existing diffusion models on graphs can be classified into two main categories depending on the type of noise injected, i.e. Gaussian or discrete. Previous works employed Gaussian diffusion models both on general graphs (niu2020permutation, ; jo2022score, ) and molecules (shi2021learning, ; xu2022geodiff, ). However, adding Gaussian noise into the adjacency matrix will destroy the sparsity of the graph, which hinders the scalability of the diffusion models (haefeli2022diffusion, ). Recent works adapted discrete diffusion models to graphs with categorical transition kernels (vignac2023digress, ; chen2023efficient, ; chen2023dexplainer, ). We denote the adjacency matrix of a graph as $\mathbf{A}^{0}\in\{0,1\}^{n\times n}$ with $n$ nodes. With details in Appendix A, we write the forward process to corrupt the adjacency matrix into a sequence of latent variables as Bernoulli distribution
$q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\alpha^{t}\mathbf{A}^{t-1}+\left(1-\alpha^{t}\right)\pi\right),$ (1)
$q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\mathbf{A}^{0}\right)=\frac{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{0}\right)}{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)},$
@@
$p_{\theta}\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t}\right)\propto\sum_{\widetilde{\mathbf{A}}_{0}}q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\widetilde{\mathbf{A}}_{0}\right)\tilde{p}_{\theta}\left(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t}\right),$ (2)
where $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t})$ represents the denoising network that predicts the original adjacency matrix from the noisy adjacency matrix. The parameters are estimated by optimizing the variational lower bound on the negative log-likelihood (austin2021structured, )
$L_{\mathrm{vb}}=$ $\sum_{t=2}^{T}\mathbb{E}_{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\mathbf{A}^{0}\right)\|\ p_{\theta}\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t}\right)\right)\right]$ (3)
$-\mathbb{E}_{q\left(\mathbf{A}^{1}\mid\mathbf{A}^{0}\right)}\left[\log p_{\theta}\left(\mathbf{A}^{0}\mid\mathbf{A}^{1}\right)\right]+\mathbb{E}_{q\left(\mathbf{A}^{0}\right)}\left[D_{\mathrm{KL}}\left(q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)\|\ p\left(\mathbf{A}^{t}\right)\right)\right].$
@@ 3 Method
In this section, our goal is to build UniAug to understand the diverse structure patterns of graphs and perform data augmentation with a range of objectives. As illustrated in Fig.1, UniAug consists of two main components: a pre-trained diffusion model and the downstream adaptation through structure augmentation. We first collect thousands of graphs from varied domains with diverse patterns. To construct a general model free of downstream inductive biases, we train a self-conditioned discrete diffusion model on graph structures. In the downstream stage, we train an MLP guidance head on top of the diffusion model with objectives across different levels of granularity. We then augment the downstream dataset by generating synthetic structures through guided generation, where the augmented graph is composed of generated structures and original node features. Subsequently, we apply the augmented data to train a task-specific model for performing downstream tasks. Below, we elaborate on the data collection process, the architecture of the discrete diffusion model, and the guidance objectives.
@@ 3.1 Pre-training data collection
Figure 2: Normalized structural properties of Network Repository and Github Star. We enlarge the distribution coverage of our collection by combining both datasets.
In light of the data scaling spirit, we expect our pre-training data to contain diverse data patterns with sufficient volume. As graphs from different domains exhibit different patterns (milo2002network, ), we wish to build a collection of graphs from numerous domains to enable a universal graph pattern library with pre-training. Within the publicly available graph databases, Network Repository (rossi2015network, ) provides a comprehensive collection of graphs with varied scales from different domains, such as biological networks, chemical networks, social networks, and many more. Among the thousands of graphs in the Network Repository, some of them contain irregular patterns, including multiple levels of edges, extremely high density, et cetera. To ensure the quality of the graphs, we analyze the graph properties following Xu et al. (xu2023better, ) and filter out the outliers. In addition, we observe that the coverage of graphs in the Network Repository is incomplete according to the network entropy and scale-free exponent, as we observe a relatively scattered space in the middle of Fig. 2. To fill in the gap, we include a subset of the GitHub Star dataset (rozemberczki2020karate, ) by random sampling $1000$ graphs into our graph collection. The selected graphs are utilized to train a discrete diffusion model.
@@ 3.2 Pre-training through diffusion model
Diffusion models have demonstrated the ability to facilitate transferability from a data augmentation perspective on the images (trabucco2024effective, ; you2024diffusion, ; he2023is, ). Unlike the traditional hand-crafted data augmentation methods, diffusion models can produce more diverse patterns with high quality trabucco2024effective . With the aid of diffusion guidance (ho2022classifier, ; dhariwal2021diffusion, ), these methods can achieve domain customization tailored to specific semantic spaces (you2024diffusion, ; he2023is, ). Despite the success of data augmentation through diffusion models on images, the non-Euclidean nature of graph structures poses challenges for data-centric learning on graphs. In addition, the fact that most graphs in the Network Repository are unlabeled exacerbates the challenges, as the absence of labeled data results in substantially lower generation quality for diffusion models (dhariwal2021diffusion, ; bao2022conditional, ).
@@
To address the aforementioned challenges, we propose to construct a self-conditioned discrete diffusion model on graph structures. Unlike Gaussian-based diffusion models, discrete diffusion models (hoogeboom2021argmax, ; austin2021structured, ; campbell2022continuous, ; vignac2023digress, ) operate with discrete transition kernels between latent variables, as shown in Section 2. The key reason we opt for the discrete diffusion models lies in the sparse nature of graphs, where adding Gaussian noise into the adjacency matrix will result in a dense graph (haefeli2022diffusion, ). On the contrary, discrete diffusion models effectively preserve the sparse structure of graphs during the diffusion process, thus maintaining the efficiency of the models on graphs.
@@
To accommodate for unlabeled graphs, we adopt a self-supervised labeling strategy as an auxiliary conditioning procedure (gao2022large, ; hu2023self, ). By leveraging the self-labeling technique, we are able to upscale the diffusion model to data with more diverse patterns (gao2022large, ). The self-labeling technique requires two components: a feature extractor and a self-supervised annotator.
@@
Feature extractor. We extract graph-level features by calculating graph properties, including the number of nodes, density, network entropy, average degree, degree variance, and scale-free exponent following xu2023better . The first two represent the scale of the graph corresponding to nodes and edges, and the rest indicate the amount of information contained within a graph (xu2023better, ). We compute the properties of one graph and concatenate them to get a graph-level representation.
@@
Self-supervised annotator. To assign labels to graphs in a self-supervised manner, we employ clustering algorithms on the graph-level representations. The number of clusters is determined jointly by the silhouette score (rousseeuw1987silhouettes, ) and the separation of the graphs. The candidates of the number of clusters are chosen to ensure different clusters are well separated. Among the candidates, we select the final number of clusters by maximizing the mean Silhouette Coefficient of all samples.
@@
Next we detail the parameterization of the denoising model $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{A}^{t})$ with the self-assigned graph-level labels $\mathbf{k}$. The denoising model recovers the edges of the original adjacency matrix by predicting the connectivity of the upper triangle, which can be formulated as a link prediction problem (zhang2018link, ; kumar2020link, ). Following the link prediction setup, the denoising model is composed of a graph transformer (GT) (shi2020masked, ) and an MLP link predictor. Denote the hidden dimension as $d$, we treat the node degrees as node features and utilize a linear mapping $f_{d}:\mathbb{R}\mapsto\mathbb{R}^{d}$ to match the dimension. Similarly, we utilize another linear mapping $f_{t}:\mathbb{R}\mapsto\mathbb{R}^{d}$ for timestep $t$ and learnable embeddings $f_{k}:\{0,\dots,K\}\mapsto\mathbb{R}^{d}$ for labels $\mathbf{k}$, where $K$ is the number of clusters. The outputs are summed together and then fed into the GT. Mathematically, we have
@@
3.3 Downstream adaptation through data augmentation
The downstream phase of UniAug is to augment the graph topology through guided generation. This guidance process serves to provide downstream semantics for the diffusion model, thus bridging the gap between the pre-training distribution and the downstream datasets. Among the techniques for diffusion guidance, gradient-based methods (dhariwal2021diffusion, ; gruver2024protein, ) offer versatile approaches by incorporating external conditions that are not present during training. For the discrete diffusion process, we opt for the gradient-based NOS method (gruver2024protein, ) due to its flexibility and efficiency. Specifically, we build an MLP regression head $g_{\theta}:\mathbb{R}^{d}\mapsto\mathbb{R}^{r}$ that takes the hidden representations $\mathbf{h}^{t}$ as the input and outputs the guidance objective of dimension $r$. Denote $\tau$ as the temperature, $\gamma$ as the step-size, $\lambda$ as the regularization strength, and $\varepsilon$ drawn from $\mathcal{N}(0,I)$, we sample from $\tilde{p}^{\prime}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\propto\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\exp\left(g_{\theta}\left(\mathbf{h}^{t}\right)\right)$ via Langevin dynamics
Node level. Node labels provide the supervision signal for node classification tasks. Beyond node labels, node degrees are a fundamental factor in the evolutionary process of a graph (liu2011controllability, ). From the perspective of network analysis, centrality measures indicate the importance of nodes from various viewpoints (borgatti2005centrality, ). Empirically, we observe that utilizing different node-level heuristics as guidance targets tends to yield similar outcomes. Therefore, we focus on node labels and node degrees.
Edge level. Edge-level heuristics can be broadly classified into two categories: local structural heuristics, such as Common Neighbor and Adamic Adar (adamic2003friends, ), and global structural heuristics, such as Katz (katz1953new, ) and SimRank (jeh2002simrank, ). Similar to node-level heuristics, empirical observations suggest that different edge-level heuristics tend to yield comparable guidance effects. In this work, we focus on the Common Neighbors (CN) heuristic due to its efficiency. Another edge-level guidance objective is to recover the adjacency matrix from the node representations in a link prediction way, similar to how we parameterize the denoising network. We anticipate that such link prediction objective helps to align the generated graph with the downstream data on the granularity of edges.
Graph level. Graph labels offer the supervision signal for graph classes or regression targets. In addition, we incorporate graph-level properties (xu2023better, ) as quantitative measures to bridge the gap between the pre-training distribution and the downstream dataset. We empirically observe that graph label guidance offers significantly higher performance boosts compared to properties on graph-level tasks. Therefore, we focus on graph labels in our experiments.
We provide our choice of objectives for each task in Appendix B. We note that all the above objectives are natural choices inspired by heuristics and downstream tasks. There exist many other self-supervised objectives to be explored, such as community-level spectral change (tan2024community, ) and motif occurrence prediction (rong2020self, ). We leave the study of objectives as one future work. With the diffusion guidance, we assemble the augmented graphs with generated structures and original node features. The augmented graphs are then fed into downstream-specific GNNs.
3.4 Comparison to existing methods